Write the nuclear equation for the nuclear reaction in which a nuclei undergoes ßdecay and changes to cesium-133. Solution: a). 𝑍? 𝐴 →−1ß 0 + 𝐶𝑠 55 133


1
Expert's answer
2021-03-15T09:10:32-0400

Solution:

A nuclei undergoes β-decay as shown below:



A = 133 + 0 = 133

Z = 55 + (-1) = 54

Hence,

133-Xe - xenon


Answer: xenon-133 (Xe) undergoes β-decay and changes to cesium-133 (Cs).